OpenStudy (jiteshmeghwal9):

The sum of the roots of \[{1 \over {x+a}}+{1 \over {x+b}}={1 \over c}\] is Zero. The products of the roots is:-

OpenStudy (jiteshmeghwal9):

\[Ans.-{{-1\over2}(a^2+b^2)}\]

OpenStudy (anonymous):

I think firstly you should solve to make it quadratic equation..

OpenStudy (jiteshmeghwal9):

I wants full solution.

OpenStudy (anonymous):

Take the LCM of the denominator...

OpenStudy (anonymous):

What is the LCM or LCD of (x+a) and (x+b) ??

OpenStudy (jiteshmeghwal9):

I gt\[{(x+b)+(x+a)}\over{(x+a)(x+b)}\]

OpenStudy (anonymous):

Yes you are right.. Now cross multiply...

OpenStudy (jiteshmeghwal9):

\[x^2+xb+ax+ab\]

OpenStudy (anonymous):

You have denominator too.. Now cross multiply: \[\frac{a}{b} = \frac{c}{d}\] \[a \times d = b \times c\]

OpenStudy (jiteshmeghwal9):

\[c(x+a+x+b)=(x^2+xb+ax+ab)\]

OpenStudy (jiteshmeghwal9):

M i right???

OpenStudy (anonymous):

Distribute c inside the parenthesis..

OpenStudy (anonymous):

Have faith on yourself.. You are going right..

OpenStudy (jiteshmeghwal9):

\[2cx+ca+cb=x^2+xb+ax+ab\]

OpenStudy (anonymous):

Now can you take all the terms to one side only.. take all the Left hand side terms to right hand side.. Remember their sign will get reverse..

OpenStudy (jiteshmeghwal9):

\[2cx+ca+cb-x^2-xb-ax-ab=0\]

OpenStudy (anonymous):

You done opposite to that.. Does not matter.. Multiply (-1) both the sides..

OpenStudy (jiteshmeghwal9):

\[-1(2cx+ca+cb-x^2-xb-ax-ab)=-1(0)\]

OpenStudy (anonymous):

Apply -1 to the brackets..

OpenStudy (jiteshmeghwal9):

\[-2cx-ca-cb+x^2+xb+ax+ab=0\]

OpenStudy (anonymous):

Now from all the terms containing x take x common.. Able to do that??

OpenStudy (jiteshmeghwal9):

I could'nt understand wht u wants to say:(

OpenStudy (anonymous):

What are the x terms there? Tell me the entire terms that contain x.

OpenStudy (jiteshmeghwal9):

-2cx,x^2,xb,ax

OpenStudy (anonymous):

I am saying the terms which contains x only not x squared and not constants.. tell me the terms having x..

OpenStudy (jiteshmeghwal9):

2cx, xb ,ax

OpenStudy (anonymous):

Now in these terms x is common and present in all the terms.. Take x common from it.. Getting??

OpenStudy (jiteshmeghwal9):

x(2c),x(b),x(a)

OpenStudy (anonymous):

From all simultaneously.. Like if I have: x + 7xb + 8xa = x(1+ 7b + 8a) Like this..

mathslover (mathslover):

Is the question answered or should i enter in between ?

OpenStudy (anonymous):

Ya sure you can give your opinion too @mathslover

mathslover (mathslover):

Ok i will like to give wait

mathslover (mathslover):

Please zoom to see it as a good file :) .. i got this and m solving further

mathslover (mathslover):

@waterineyes any suggestion ?

OpenStudy (anonymous):

Very Good..

mathslover (mathslover):

But what next ?

mathslover (mathslover):

I am just going all around :

OpenStudy (anonymous):

the equation looks like x^2+(a+b-2c)x +ab-ac-bc=0 but given a+b-2c=0 i.e. a+b=2c product of roots is ab-ac-bc=ab-c(a+b) =ab-((a+b)^)/2= 1/2 (2ab -a^2 -b^2 -2ab) =-1/2 (a^2 +b^2)....ans

OpenStudy (anonymous):

I got it.. Wait..

OpenStudy (jiteshmeghwal9):

i also gt thnx a lot & here the question closes;)

OpenStudy (anonymous):

mention not..

OpenStudy (anonymous):

\[x^2 + bx + ax + ab - ac - bc - 2cx = 0\] \[x^2 + (a+b-2c)x + (ab -ac-bc) = 0\] From here, A = 1, B = (a + b - 2c) and C = (ab -ac- bc) Sum of Roots = 0 \[-(a+b -2c) = 0\] \[c = \frac{a+b}{2}\] Product of Roots = C \[ab - ac - bc = ab - a(\frac{a+b}{2}) - b(\frac{a+b}{2})\] \[= \frac{2ab - a^2 - ab - ba - b^2}{2} = \frac{-a^2 - b^2}{2} = \frac{-1}{2}(a^2 + b^2)\]

mathslover (mathslover):

good work @matricked and @waterineyes

OpenStudy (anonymous):

Thanks buddy..

OpenStudy (anonymous):

@mathslover mention not